I have two questions.
1) Write a quadratic equation with the variable x having the given numbers. Write an equation in standard form ax^2+bx+c=0 solutions are -3,1
2) There is a cookbook open to two pages the product of the page numbers are 1190. Find the page numbers.
first page-
second page-
Thanks in advance

Answer by stanbon(75887)

1) Write a quadratic equation with the variable x having the given numbers. Write an equation in standard form ax^2+bx+c=0 solutions are -3,1
---
f(x) = (x+3)(x-1)
f(x) = x^2 + 2x -3
--------------------------------------------------------
2) There is a cookbook open to two pages the product of the page numbers are 1190. Find the page numbers.9
---
The page numbers are consecutive natural numbers:
1st : x
2nd : x+1
-----
Equation:
x(x+1) = 1190
x^2 + x - 1190 = 0
(x+35)(x-34) = 0
x = 34 (first page
x+1 = 35 (next page
=============================
Cheers,
Stan H.